ho:u1-u2=2 and h1:u1-u2=2


1
Expert's answer
2021-06-17T15:44:42-0400

h0: u1 - u2 = 2

u1 = u2 + 2

It means u1 > u2

Therefore, mean of group u1 > .mean of group u2.

And h0 will be called alternative.


h1: u1 - u2 = 2

u1 = u2 + 2

It means u1 > u2

Therefore, mean of group u1 > mean of group u2.

And h1 will be called alternative.